OpenStudy (anonymous):

If f(x) = −4x + 1, find f(−5).

hartnn (hartnn):

just put x=-5 so, what is -4(-5)+1 =... ?

Parth (parthkohli):

@ahoward79: You have asked many similar questions. Did you get the concept? =)

hartnn (hartnn):

O.o

OpenStudy (anonymous):

@ParthKohli No :/ I just dont understand. Do you think you could work one all the way out so I can see how to do it?

OpenStudy (anonymous):

connections academy?

OpenStudy (anonymous):

You need to plug in -5. so here it's : -4(-5) + 1 which is 21.

OpenStudy (hba):

It's just a plug in game @ahoward79

OpenStudy (hba):

Just replace 'x' with '-5'

hartnn (hartnn):

if you have to find f(a) when f(x) is given, you just replace 'x' in f(x) by a. that is put x=a example : if f(x) = 5x-4 f(2) = 5(2)-4 = 10-4 =6

OpenStudy (anonymous):



OpenStudy (anonymous):

So would the answer be 21?

OpenStudy (hba):

Yeah :)

hartnn (hartnn):

but did you get how ?

OpenStudy (anonymous):

Yes :) Thank you all! I wish I could give you all medals :)

OpenStudy (hba):

@ahoward79 Do you really understand this ?

hartnn (hartnn):

so if i give you same type of other problem, will you be able to solve it ?

OpenStudy (hba):

For example : f(x)=2x-3 What will f(-1)=..

OpenStudy (anonymous):

-5

OpenStudy (hba):

Good going (y)

OpenStudy (hba):

Please show the steps how you did it :)

OpenStudy (anonymous):

2(-1) - 3

OpenStudy (hba):

Yes continue...